# Incorporating Gradients to Rules: Towards Lightweight, Adaptive Provenance-based Intrusion Detection

### v1

@@ VI-C2 Runtime Overhead
Another important metric reflecting the efficiency of a PIDS is the runtime overhead to the system. In this section, we evaluated the runtime overhead of Captain during detection and conducted a comparative study with other SOTA PIDS. We also analyzed the memory consumption of Captain during the training stage.
Runtime Overhead in Detection: We used the Python resource module to evaluate detectors’ resource consumption. We ran the detectors on CPU mode. As admitted in [51], GPUs may not be available in most real-life threat detection scenarios. We calculated the total CPU time in user mode to evaluate the CPU resource consumption. The result is shown in Fig. 4(a). Please note that the total CPU time used in Fig. 4(a) could be more than the detection time in Table VI because of the multi-core usage of CPU in the deep learning module. Since Captain only relies on straightforward arithmetic operations on real-number tags, it requires significantly less CPU time than the embedding-based PIDS (around 2% of Flash and around 0.02% of Kairos), which need complex matrix and vector operations due to their use of neural networks.
We used the psutil library in Python to monitor the live memory usage of the detectors. The comparison of memory usage over time for detection is shown in Fig 4(b). Compared to the baselines, Captain finishes detection much faster than the baselines (as we have already seen in Table VI) and achieves the lowest memory usage throughout the entire detection process.
@@
Figure 4: Comparison of resource consumption when detecting on DARPA Engagement 3 Cadets
Runtime Overhead in Training: In real-world scenarios, the training stage of the detection system can be performed offline, where users can allocate ample time and computing resources, including GPUs. Therefore, we did not evaluate the runtime overhead and the time used for training. However, as Captain records and propagates the gradients for each tag among the big graph, a reasonable concern is that saving those gradients would consume much memory when there is a dependency explosion problem. Our evaluation shows that, in practice, we would not need so much space because most gradients would always be zero since the provenance graph is a sparse graph. We evaluated the number of non-zero gradients that Captain needs to save during training. The results, shown in Fig. 8, illustrate that most nodes only need to store a small number of gradients during training. Due to the space limit, the detailed results can be found in § C-B.
@@ VI-D Resilience Against Adversarial Attacks
As more PIDS have been developed to analyze provenance data, adversarial attacks aimed at evading these provenance-based detection mechanisms have garnered more attention. In this section, we discuss the robustness of Captain against two main-stream adversarial attacks: adversarial mimicry attack and dataset poisoning attack.
@@ VI-D1 Adversarial Mimicry Attack
Mimicry attacks on PIDS involve altering the provenance data and incorporating more benign features to ”mimic benign behavior”, thereby evading detection. In this section, we evaluated the robustness of Captain against mimicry attacks using the attack methodology in [15, 34], i.e. inserting benign structures into the attack graphs. Our mimicry attack contains two steps. First, we extract some events of benign system entities from the normal training data. Then we add some ”fake” events where we replace the benign system entities with the attack entity to simulate the attack entity doing similar activities as the benign entities. To verify the effectiveness of our mimicry attack, we also use Flash as the baseline.
For evaluation purposes, we used the Cadets dataset in Engagement 3 as the example because of its relatively small scale. The details about our mimicry experiment can be found in the Appendix C-C. The result is shown in Fig. 5.
@@
In general, both Captain and Flash showed robustness against our mimicry attack attempt to some extent. The attack entity /tmp/test is still detected by both systems even after we add some normal activities. However, we can see a significant drop in the anomalous score of Flash when we introduce a relatively small number of events (fewer than 10), which is also confirmed in their paper [34]. On the other hand, Captain remains unaffected by the mimicry attack, demonstrating superior robustness compared to the baseline.
@@
The robustness of Captain results from three main reasons. First of all, the propagation of Captain is guided by heuristic-based rules, while the propagation and aggregation of Flash and other embedding-based PIDS are based on the neural network. Second, Captain is an event-level detector with a finer granularity than entity-level detectors. As mentioned before, Captain gives the detection result for each streaming event without any buffer time, which means any mimicry insertion after the real attack activities is useless. For those PIDS that need a buffer time, any events within the same time window or related to the same entity can affect the final detection result. Third, Captain is not an anomaly-based detector. In other words, the detection is performed based on the conservative, heuristic rules, while the normal data is only used to reduce the false alarms. Therefore, adding normal features can not compromise the detection capability as it is guaranteed by the conservative detection rules.
VI-D2 Traning Set Poisoning Attack
@@
Dataset poisoning refers to the malicious manipulation on the training set that is relied on by the learning system. In the context of cyber security and intrusion detection, it usually involves polluting the benign training set with adversarial events/entities. When the detector is trained on a ”benign” dataset that has been polluted, it learns patterns or features introduced by attackers. These patterns and features, once learned, could assist attackers in evading detection in future instances.
We add a regularizer term to the loss function to control the sensitivity of Captain to the training set. This enhances Captain’s robustness against dataset poisoning attacks.
@@
We evaluated this robustness using a specific real example of Pine Backdoor & Phishing Email attack from T-3. In this attack, a vulnerable pine connected to the C2 server 128.55.12.73 and executed an email attachment. However, IP 128.55.12.73 had multiple activities and caused some “false” alarms in the training set. Please note that this violates our assumption stated in § II-C that the training set must not be compromised by malicious entities or events. Nevertheless, this makes it a perfect example to test the robustness against data poisoning attacks.
Table VII shows the values of two critical parameters in detecting this attack: the initial integrity tag of 128.55.12.73 and the tag propagation rate of event (pine, read, 128.55.12.73), under different learning settings. We use the conservative setting ($A_{0}$, $G_{0}$, and $T_{0}$) as the baseline. The result illustrates that if we don’t use the regularization term, 128.55.12.73 will be treated as a pure benign IP, and the propagation rate to the process pine is 0, which lets us miss the attack. With the increase of the regularizer coefficients, Captain is more cautious when learning the training data. Finally, we successfully captured this special attack with a larger regularizer coefficient. In practice, the regularizer coefficients can be set based on the confidence in the training data, the tolerance to the missing alarm, and the trade-off between TP and FP.
TABLE VII: Case study on the effectiveness of the regularization term in the loss function.
Base The value of $\alpha,\lambda,\tau$
00
$10$
$100$
$a_{(128.55.12.73)}$ 0 1.00 0.818 0.391
$g_{(pine,read,128.55.12.73)}$ 1 0.980 0.985 0.992
Detected? Yes No No Yes
@@ VI-E1 Individual Adaptive Parameters
We conduct an ablation study on each adaptive parameter, comparing the detection performance with or without specific adaptive parameters. As each parameter can be learned and adjusted independently, we have seven separate experiments, each optimizing a different subset of the parameters, i.e. $\{A\}$, $\{G\}$, $\{T\}$, $\{A,G\}$, $\{A,T\}$, $\{G,T\}$, and $\{A,T,G\}$.
According to the experiments, adjusting the threshold ($T$) alone may not have much effect on the detection results because when the initial tags and the propagation rates are unchanged, the tags after the propagation also tend to be 0 or 1. Therefore, tuning a continuous threshold would not influence the final alarm decision. It should be noted that tuning $A$ plays the most major role in mitigating false alarms among these three adaptive parameters. In other words, tuning the initial tags of nodes can help filter out the false alarms more than tuning the propagation rates and the thresholds.
While tuning parameters $A$, $G$, and $T$ together often yields incremental improvements, adjusting $A$ alone sometimes results in fewer false alarms in certain datasets (S-3 and C-3). This is because, without the benefit of adjusting the propagation rate and the threshold, we have to assign higher integrity tags to more nodes to counteract false alarms in training sets. Although this can reduce false alarms more crudely and simply, it also carries the risk of being excessively lenient, which might be hazardous in some contexts.
Another advantage of tuning all these three parameters is that it can accelerate the training process. This is shown in Fig.6. Although $\{A\}$, $\{A,T\}$, and $\{A,T,G\}$ achieve similar good detection results on the testing set, $\{A,T,G\}$ converges using fewer epochs during training.
@@ VI-E2 Learning Rate in Optimization
During training, the learning rate is a critical hyperparameter that not only determines the learning speed but also controls the granularity of the search. We conducted an experiment on different learning rates on the DARPA Engagement 3 Cadets data. The detailed results can be found in § C-D. As illustrated in Fig. 9, a higher learning rate usually leads to a sharper loss decline during the initial training epochs. However, for ongoing training, a lower learning rate may yield a finer-tuned search process and lower loss. In cases where training time on large graphs is a critical factor, a balance between training speed and training quality may be necessary.
@@ VI-F Case Study
To explore the explainability of Captain, we delve into the training outcomes of the E3-Trace dataset as a case study.
Tag Initialization (A) Adjustments: Examining the parameters in $A$, the adjustments from $A_{0}$ primarily focused on network objects, showcasing a broad spectrum of modifications. Notably, internal IP addresses, such as 128.55.12.10, 128.55.12.122 and 128.55.12.73, were assigned initial integrity values close to 1. Other IP addresses, like 65.214.39.18 and 216.9.245.101, are assigned initial integrity values around $0.5$. Additionally, certain IP addresses, such as 208.17.90.11 and 162.93.202.80, were endowed with initial integrity values slightly exceeding 0. The discrepancies in the learned integrity values reflect the varying levels of trust Captain places in these IP addresses. Internal IP addresses, in particular, triggered numerous alarms through potentially malicious behaviors during the training period. To reduce such false alarms, Captain learned to trust these internal IP addresses, leading to the assignment of elevated integrity values. Other IP addresses, appearing less frequently and engaging in less sensitive activities, received neutral integrity values. Moreover, some IP addresses, despite participating in sporadic sensitive activities eventually deemed benign, did not garner significant trust from Captain. Consequently, their initial integrity values underwent a minimal change from $A_{0}$. Compared with manual whitelisting, Captain can assign integrity with finer granularity, allowing it to navigate the trade-offs between true positives and false positives more effectively.
Propagation Rate (G) Adjustments: Examining the parameters in $G$, the adjustments from $G_{0}$ predominantly focus on read and write operations on files. Notably, events involving bash reading from .bash_history received a relatively low propagation rate, approximately $0.5$, while other events associated with .bash_history maintained a high propagation rate close to 1. As discussed in § III, bash frequently engages in reading and writing activities with .bash_history during routine tasks, potentially causing dependency explosions and triggering numerous false alarms. Through the training process, Captain learned to mitigate false positives by reducing the propagation rates for events involving bash reading from .bash_history. Meanwhile, Captain remained cautious with other .bash_history related events, maintaining relatively high propagation rates to ensure malicious intent can still propagate through alternative pathways. The fine-grained adjustment of propagation rates by Captain diminishes false positives while preserving its robustness, thereby preventing attackers from easily evading detection.
Alarm Generation Threshold (T) adjustments: Upon scrutinizing the parameters in $T$, the adjustments encompass various events capable of triggering alarms. Notably, we observed that the threshold for generating data leak alarms is set close to 0 for processes such as sshd, thunderbird, and firefox when they send messages to internal IP addresses. In contrast, the threshold for generating data leak alarms when these processes send messages to other IP addresses remains around the default value of $0.5$. Furthermore, the internal IP addresses associated with these processes are distinct. For instance, sshd exhibits a low alarm generation threshold exclusively when sending messages to 128.55.12.122 and 128.55.12.10, while thunderbird displays a low threshold when communicating with 128.55.12.73. These observations indicate that Captain has effectively learned to diminish false alarms by lowering the threshold for generating such alarms at the edge level. These adjustments are finely tuned to specific processes communicating with particular IP addresses, while leaving the alarm generation threshold unaltered for other processes communicating with these IP addresses or for these processes communicating with other IP addresses. This fine-grained and nuanced approach not only allows Captain to effectively reduce false alarms but also fortifies it against living-off-the-land attacks that seek to emulate benign activities. The resilience to such attacks persists unless the attackers can successfully gain access to the internal IP addresses featured in the training dataset, thereby significantly heightening the complexity and challenge of executing such an attack.
@@ VII Discussion and Future Work
VII-A More Optimization Algorithms
In this paper, we utilize the gradient descent algorithm for optimization because of its simplicity and clearness. There are more gradient-based optimization algorithms. We discuss their applicability based on our differentiable tag propagation framework in this section. In fact, we used Batch gradient descent in Captain. The batch size is fixed to the number of events in each dataset. Using a smaller batch size means updating the adaptive parameters before finishing a single round of detection, which could cause tag inconsistency in a single provenance graph. Instead of directly updating parameters based on the gradients, we can modify these gradients to reach optimal values more quickly and accurately, using algorithms such as Nesterov Accelerated Gradient (NAG) [31], Adadelta [49], RMSprop, Adagrad [14], Adam [21], AdaMax, Nadam, and so on. These algorithms require the first-order gradient, which is offered by the differentiable tag-propagation framework in Captain. We leave the implementation of these algorithms as the extensions of Captain.
VII-B Online Learning
We assess Captain in an offline setting for automatic data labeling. However, the architecture of the detection and learning modules doesn’t preclude the use of Captain in an online setting with manual labeling. Specifically, Captain’s learning module can compute loss upon receiving manually determined false alarms. Subsequently, the learning module can promptly update the adaptive parameters for subsequent detection.

### v2

@@ VI-C2 Runtime Overhead
Another important metric reflecting the efficiency of a PIDS is the runtime overhead to the system. In this section, we evaluate the runtime overhead of Captain during detection and conduct a comparative study with other SOTA PIDS. We also analyze the memory consumption of Captain during the training stage.
Runtime Overhead in Detection: We use the Python resource module to evaluate the resource consumption of detectors on CPU mode. As admitted in [13], GPUs may not be available in most real-life threat detection scenarios. We calculate the total CPU time in user mode. The result is shown in Fig. 4(a). Please note that the total CPU time used in Fig. 4(a) could be more than the detection time in Table VI because of the multi-core CPU usage by deep learning models. Since Captain only relies on straightforward arithmetic operations on real-number tags, it requires significantly less CPU time than the embedding-based PIDS (around 2% of Flash and around 0.02% of Kairos), which need complex matrix and vector operations due to their use of neural networks. Moreover, Captain’s fine-grained detection rules only increase CPU time by a modest 5.6% compared to Morse. Given the significant improvement on Morse in detection accuracy (over 90% reduction in false alarms) and the substantial decrease in CPU time compared to other embedding-based PIDS, we believe the slightly higher CPU usage than Morse is acceptable in real-world detection scenarios.
We use the psutil library in Python to monitor the live memory usage during detection. The comparison of memory usage over time is shown in Fig 4(b). Captain finishes detection much faster than the embedding-based PIDS (as already shown in Table VI) and achieves the lowest memory usage throughout the entire detection process. The memory usage and detection time of Captain and Morse are similar, shown by the overlapping curves in Fig. 4(b). The slight difference in memory usage between Captain and Morse is due to the storage of fine-grained rules. In general, Captain retains the lightweight and fast characteristics of the rule-based PIDS.
@@
Figure 4: Comparison of resource consumption when detecting on DARPA Engagement 3 Cadets. In Fig. 4(b), the memory usage curve of Captain and Morse are overlapped, showing their similar efficiency performance.
Runtime Overhead in Training: The training of the detectors can be performed offline, where users can allocate ample time and computing resources, including GPUs. Therefore, we did not evaluate the runtime overhead and the time used for training. However, as Captain records and propagates the gradients for each tag among the big provenance graph, a reasonable concern is that saving those gradients would consume much memory when there is a dependency explosion problem. We evaluated the number of non-zero gradients saved by Captain during training. The results, shown in Fig. 8, illustrate that most nodes only need to store a small number of non-zero gradients during training. Due to the space limit, the detailed analysis can be found in §E-A.
@@ VI-D Resilience Against Adversarial Attacks
In this section, we discuss the robustness of Captain against two main-stream adversarial attacks: adversarial mimicry attack and training set poisoning attack.
@@ VI-D1 Adversarial Mimicry Attack
Mimicry attacks on PIDS involve altering the provenance data and incorporating more benign features to “mimic benign behaviors”, thereby evading detection. In this section, we evaluated the robustness of Captain against mimicry attacks using the attack methodology in [35, 16], i.e. inserting benign structures into the attack graphs. Our mimicry attack contains two steps. First, we extract some events of benign system entities from the normal training data. Next, we create “fake” events by substituting the benign entities with the attack entity, simulating the attack entity performing activities similar to those of the benign entities. To verify the effectiveness of the mimicry attack, we use Flash as the baseline. For evaluation purposes, we use the Cadets from Engagement 3 because of its relatively small scale. The details about our mimicry experiment can be found in the Appendix E-B. The result is shown in Fig. 5.
@@
In general, both Captain and Flash showed robustness against our mimicry attack attempt to some extent. The attack entity /tmp/test is still detected by both systems even after we add some normal activities. However, we see a significant drop in the anomalous score of Flash when we introduce a relatively small number of events (fewer than 10), which is also confirmed in their paper [16]. On the other hand, Captain remains unaffected against the mimicry attack, demonstrating superior robustness compared to the baseline.
@@
The robustness of Captain results from three main reasons. First of all, the propagation of Captain is guided by heuristic-based rules, while the propagation and aggregation of Flash and other embedding-based PIDS are based on the neural network. Second, Captain is an event-level detector with a finer granularity than entity-level detectors. As mentioned before, Captain gives the detection result for each streaming event without any buffer time, which means any mimicry insertion after the real attack activities is useless. For those PIDS with the buffer time, any events within the same time window and related to the same entity can affect the final detection result. Third, Captain is not an anomaly-based detector. In other words, the detection is performed based on the conservative, heuristic rules, while the normal data is only used to reduce the false alarms. Therefore, adding normal features can not compromise the detection capability as it is guaranteed by the conservative detection rules.
VI-D2 Training Set Poisoning Attack
@@
Training set poisoning refers to the malicious manipulation of the training set. In the context of cyber security and intrusion detection, it usually involves polluting the benign training set with adversarial events/entities. When the detector is trained on a “benign” dataset that has been polluted, it learns patterns or features introduced by attackers. These patterns and features, once learned, could assist attackers in evading detection in future instances.
@@
We add the regularizer term to the loss function to control the sensitivity of Captain to the training set. This enhances Captain’s robustness against dataset poisoning attacks. In this section, we present a real case of Pine Backdoor & Phishing Email attack from Engagement 3 Trace. In this attack, a vulnerable pine connected to the email server 128.55.12.73, downloaded and executed an email attachment. However, IP 128.55.12.73 had multiple activities and caused some “false” alarms in the training set. Please note that this violates our assumption stated in § II-C that the training set must not be compromised by attack entities or events. Nevertheless, this makes it a perfect example to test the robustness against data poisoning attacks.
We analyze the critical parameters for detecting this attack: the initial integrity tag ($a$) of IP node 128.55.12.73, to illustrate the effect of the regularization term. Without the regularization term in the loss function, 128.55.12.73 will be assigned a high integrity score (near 1.0) due to its activities in the training set, which lets us miss the phishing email attack starting from it. However, with the help of the regularizer terms ($\alpha=10$), Captain is more cautious during training. Like the “grey node” motivating example in §III, Captain assigns the integrity score as the mediocre 0.488, which removes many false alarms caused by this IP while successfully capturing the phishing email attack during testing (more details can be found in §VI-F). The propagation rate of event (pine, read, 128.55.12.73) is also more conservative against the polluted training set due to the regularizer term.
@@ VI-E1 Individual Adaptive Parameters
We conduct an ablation study on each adaptive parameter. As each parameter can be learned independently, we have seven separate experiments, each optimizing a subset of the parameters, i.e. $\{A\}$, $\{G\}$, $\{T\}$, $\{A,G\}$, $\{A,T\}$, $\{G,T\}$, and $\{A,T,G\}$.
According to Fig.6, adjusting the threshold ($T$) alone may not have much effect on the detection results because without changing the initial tags and the propagation rates, the updated tags still tend to be 0 or 1. Fig.6 also shows that tuning all these three parameters can accelerate the training process. Although $\{A,T\}$ and $\{A,T,G\}$ achieve similar good detection results on the testing set, $\{A,T,G\}$ converges using fewer epochs during training. It should be noted that tuning $A$ plays the most major role in mitigating false alarms among these three adaptive parameters. In certain datasets (S-3 and C-3), adjusting $A$ alone sometimes results in fewer false alarms than tuning parameters $A$, $G$, and $T$ together because, without the benefit of learning propagation rates and thresholds, we have to assign higher integrity scores to more nodes to counteract false alarms during training. However, reducing false alarms more crudely carries the risk of being excessively lenient.
@@ VI-E2 Learning Rate in Optimization
We conducted an experiment on different learning rates on the DARPA Engagement 3 Cadets. The detailed results can be found in § E-C. As illustrated in Fig. 9, a higher learning rate usually leads to a sharper loss decline during the initial training epochs. However, for ongoing training, a lower learning rate may yield a finer-tuned search process and lower loss, necessitating a balance between training speed and training quality.
@@ VI-F Case Study
To illustrate the interpretability of Captain’s detection process, we delve into the training outcomes of the E3-Trace dataset as a case study. These cases also show the improvement of Captain compared to existing rule-based PIDS.
VI-F1 Better Detection Capability due to Fine-Grained Rules
The most significant improvement of Captain is its fine-grained detection capability, enabled by the adaptive parameters. As mentioned in § III, unlike the discrete, human-assigned trustworthiness levels and universal rules used in [22, 21, 23], the fine-tuned numerical parameters of each entity and event in the graph can better help distinguish nuances between benign and malicious patterns.
Fig. 7(a) demonstrates how Captain outperforms the baselines in distinguishing between two similar behaviors using fine-tuned parameters. In both cases, process pine reads from 128.55.12.73 (possibly the email server) and creates, writes, and changes the permissions of a new file. It is difficult for previous rule-based PIDS [21, 22] to assign a tag for 128.55.12.73. Take the SOTA Morse as an example - if Morse initialize the node 128.55.12.73 as a benign node with the data integrity score as 1.0, it would miss at least the first stage of the attack. On the other hand, if Morse assigns the initial tag as “Untrusted” (this is what they did in order to capture all attacks), it would trigger multiple false “Malicious File Creation” and “Change Permissions” alarms on the benign graph. Instead, Captain assigns a moderate initial tag $a_{n}$, 0.488, to this IP and slightly tunes down the threshold $thr$ for the event (pine, create, /home/admin/pinerc174500) from 0.5 to 0.478. As $0.478<0.488$, the creation of the file will not be alarmed, but in real attacks, the $thr$ for the event (pine, create, /tmp/tcexec) is still 0.5, meaning creating a malicious file would still trigger alarms ($0.5>0.488$).
(a) The graph on the left side represents normal behavior in the training set, while the graph on the right side depicts an attacker executing a malicious email attachment.
(b) Captain assigns different propagation rates ($g_{e}$) to events to control dependency explosion more precisely.
Figure 7: Case study from Engagement 3 Trace. We omitted some irrelevant details to keep the graphs clean and succinct.
Captain’s fine-grained rules also help address the dependency explosion issue caused by libresolv.so.2. Unlike Morse, which sets universal decay/attenuation factors for all processes, Captain learns distinct propagation rates for different events. As shown in Fig. 7(b), Captain assigns lower propagation rates to events that frequently cause dependency explosions (e.g., being read by sshd or firefox) while maintaining high propagation rates for other events. Even for the same process, firefox, Captain learns to suppress the dependency propagation when reading this file ($g_{e}=0.63$) and remains sensitive to maliciousness propagation when loading this file ($g_{e}=0.99$). Consequently, Captain avoids being overly lenient for potential attack events when addressing the dependency explosion issue.
The fine-grained parameters of Captain also outperform many embedding-based PIDS in detecting mimicry attacks. For instance, in Fig. 7(a), the attacker can modify the file name of /tmp/tcexec with the knowledge that pine might interact with a file like /home/admin/pinerc+numbers. Since generalization techniques like Word2Vec and Sentence2Vec are widely used in embedding-based PIDS, they can manipulate the file name by appending random digits to pinerc, making it similar to pinerc174500 after embedding. However, to evade Captain, the attacker must know the exact file name of pinerc174500. In addition, deep learning usually relies on the amount of training data. But there are only four events related to file /home/admin/pinerc174500 during the 7-day training set, making it challenging for the neural networks to learn such patterns.
VI-F2 Robustness brought by Differentiable Tag-Propagation Framework
Some previous rule-based PIDS [7] can customize their detection system using benign training data. However, the adjustments are based on the training set rather than the detection results. For example, NoDoze assigns a distinct anomaly score to each event based on its frequency in the training set. However, due to training set poisoning and living-off-the-land attack techniques, the frequent events in the training set cannot be fully trusted. For instance, the events related to dbus-daemon are very common in the training set of Trace in Engagement 3. Typical events such as dbus-daemon writes or reads /var/run/dbus/system_bus_socket occur over 100,000 times during seven days. Such events and the dbus-daemon entity would be assigned with a low anomaly score in frequency-based systems like NoDoze. However, the low anomaly scores allow the attackers to deceive the detectors using living-off-the-land techniques related to the D-Bus process [56]. In contrast, Captain does not change the parameters related to dbus-daemon because, despite the large number of events associated with dbus-daemon in the training set, no false alarms are triggered by it. Captain adjusts the parameters related to an event based on the false alarms caused by it rather than the frequency of it.
@@ VII Discussion and Future Work
VII-A Overfitting in Parameter Learning
Overfitting is an important issue in machine learning. As Captain learns the benign patterns in the one-class manner starting from conservative settings, overfitting is, however, a preferred and safer strategy. In one-class learning, we want to generate a tightest boundary for the target class (benign patterns). Therefore, Captain avoids generalizing excessively to minimize the risk of evasion, as demonstrated in our example in §VI-F.
VII-B More Optimization Algorithms and Online Learning
In this paper, we utilize the gradient descent algorithm for optimization due to its simplicity and clearness. We can reach optimal values more quickly and accurately, using algorithms such as Nesterov Accelerated Gradient (NAG) [57], Adadelta [58], RMSprop, Adagrad [59], Adam [60], AdaMax, Nadam, and so on. These algorithms require the first-order gradient, offered by the differentiable detection framework in Captain. We leave the implementation of these algorithms as the extensions of Captain.
Long-term maintenance is important in real-world scenarios. Although we assess Captain in an offline setting, Captain’s learning module can compute loss upon receiving the feedback for each event. Subsequently, the learning module can promptly update the adaptive parameters for subsequent detection.